OpenStudy (anonymous):

Given f(2) = 12, which of the following could be an equation of the function? A. f(x) = 8x – 2 B. f(x) = 4x + 4 C. f(x) = 6x – 4 D. f(x) = –4x – 4

OpenStudy (anonymous):

Substitute x as 2 in every option and check it comes to 12

OpenStudy (anonymous):

f(2) = 12 means that when x = 2, the value of the function is 12. So go through each choice and see if when x = 2, the function has a value of 12.

OpenStudy (anonymous):

whats the answer @Easyaspi314

OpenStudy (anonymous):

U try it first

OpenStudy (anonymous):

i dont know this

OpenStudy (anonymous):

Let's try choice A to see if that works. f(x) = 8x - 2 Substitute x = 2, we get 8(2) - 2 = 16 - 2 = 14. No good. We want that when x = 2, f has a value of 12. So you try the remaining choices and determine which one is correct.

OpenStudy (anonymous):

C ? @Easyaspi314

OpenStudy (anonymous):

why do u think so?

OpenStudy (anonymous):

C? 6x - 4? Let's see. when x = 2, 6x - 4 = 6(2) - 4 = 12 - 4 = 8. No good.

OpenStudy (anonymous):

D ? @Easyaspi314

OpenStudy (anonymous):

Show me why you think D is correct.

OpenStudy (anonymous):

i dont knwo thats just my best answer? @Easyaspi314

OpenStudy (anonymous):

Not good enough. All what you have to do is to substitute x = 2 into that expression. Can you do that?

OpenStudy (anonymous):

is im right? @Easyaspi314

OpenStudy (anonymous):

@lad7 I assume you are here to learn, and not just to get answers. If you think D is correct, what work did you do to get choice D?